What volume does 100.0 g of sulfur trioxide occupy at STP?
1
Expert's answer
2020-09-23T04:36:25-0400

` 100 g of Sulfur Trioxide Have moles = 100 / 80 = 1.25 mol .


Volume of Sulfur Trioxide = ( 1.25 * 22.4 ) liter at STP = 28 LITER . ANSWER .